OpenStudy (kpop4life123):

There are (108)2 ⋅ 100 candies in a store. What is the total number of candies in the store?

Nnesha (nnesha):

\[\huge\rm (10^8)^2 \times 10^0\] this is your question right ?

OpenStudy (kpop4life123):

YES @Nnesha

Nnesha (nnesha):

alright so next tie use `^` for exponents

Nnesha (nnesha):

you should know two exponent rules to solve that question first one is \[\huge\rm (x^m)^n = x^{ m \times n}\](multiply both exponents ) and then 2nd one is when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]